# Dicke superradiant enhancement of the heat current in circuit QED

### v1

@@ II.1 Steady-State Density Matrix and Effective Temperature
Since we are interested in computing steady-state heat currents, we start discussing the steady-state solution of the Lindblad master equation (9). In the case of a single reservoir $\mathcal{D}_{i}(\hat{\rho})$, the steady-state solution corresponds to the thermal distribution (within each subspace with fixed $J$) characterized by the corresponding temperature $T_{i}$, since the rates satisfy the detailed balance condition
@@
with $i={\rm h},{\rm c}$, capturing the equilibrium relation between absorption and emission rates in each bath. This result, consistent with the intuition from thermodynamics, comes from using a master equation written in terms of the global jump operators for the coupled qubits (see Refs. Cattaneo et al., 2019; Chiara et al., 2018 for a discussion of this point).
In the case of two baths, the total dissipator in Eq. (10) can be expressed as:
@@
which, in turn, defines an effective temperature $T_{c}\leq T^{*}\leq T_{h}$. We can invert the previous equation to get an explicit expression for the effective temperature
[…]
Figure 3: (Color online) This figure shows the effective temperature $T^{*}$ as a function of the temperature of the hot bath, $T_{\rm h}$ (both renormalized by $T_{0}$) for various values of the cold bath, $T_{\rm c}$. Other parameters are $\gamma_{\rm h}=\gamma_{\rm c}$.
@@
In Fig. 3 we show the effective temperature $T^{*}$ as a function of the temperature of the hot bath, $T_{\rm h}$ for various values of $T_{\rm c}$. Notice that this dependence is shown only for $T_{\rm h}\geq T_{\rm c}$.
The steady-state solution $d\hat{\rho}/dt=0$, determined setting $\mathcal{D}(\hat{\rho})=0$, is thus a thermal state at temperature $T^{*}$ within each subspace with fixed $J$. However, since $\hat{J}^{2}$ is conserved, the relative occupation of different subspaces is fixed by the initial state of the qubits. Assuming that the initial state does not contain any coherence between subspaces with different $J$, the stationary state $\hat{\rho}^{(\rm s)}$ is given by
@@
where $m_{J}$ are the eigenvalues of $\hat{J}_{z}$ compatible with $J$, $P_{J}$ is the occupation probability of subspace $J$ determined by the initial state, and |$\ket{J,m_{J}}$ are the Dicke states, given in Eq. (8).
The dissipator in Eq. (14) fixes the conditional probabilities $P(m_{J}|J)$ to be thermal with temperature $T^{*}$ fixed by the detailed balance equation, Eq. (16), according to
@@
where $Z_{J,T}=\sum_{m_{J}}\exp\left[{-m_{J}T_{0}}/T\right]$ is the partition function.
@@ II.2 Heat Current
The general expression for the total heat current $\dot{Q}$ flowing out of the thermal baths can be defined as Alicki (1979); Binder et al. (2019)
@@
Using the specific forms of the Lindblad operators for the cold and hot baths, as detailed in App. C, we arrive at the following expressions for the heat currents
@@
where $i={\rm h,c}$ and $\langle\hat{x}\rangle\equiv\text{Tr}[\hat{\rho}\hat{x}]$. This result reminds the well known superradiant energy emission in an ensemble of excited qubits Raimond et al. (2001); Gross and Haroche (1982); Yan et al. (2023); Masson et al. (2023). In that context the system is typically considered to be in contact with a single cold bath at zero temperature, corresponding to setting $\gamma_{\rm h}=0$ in our case, and is initially prepared in a highly excited state. Consequently, the superradiant heat current flowing into the cold bath is given by $\dot{Q}_{\rm c}=-\hbar\omega_{0}\gamma_{\rm c}\langle\hat{J}_{+}\hat{J}_{-}\rangle$. The term $\langle\hat{J}_{+}\hat{J}_{-}\rangle$ can be expressed in terms of the individual qubit raising and lowering operators, $\hat{\sigma}_{+}^{(j)}$ and $\hat{\sigma}_{-}^{(j)}$, as $\sum_{j,j^{\prime}}\langle\hat{\sigma}^{(j)}_{+}\hat{\sigma}^{(j^{\prime})}_{-}\rangle$. This summation over collective operators encompasses $N^{2}$ terms, which could lead to a super-extensive scaling effect due to the coherence among different qubits. However, it is important to note that this superradiant effect is present only in the transient dynamics. Over time, the qubits will eventually reach thermal equilibrium with the zero-temperature bath, resulting in the cessation of the heat current.
Going back to the case under study, our primary focus is on the steady-state heat current that flows between the two baths. At this steady state, the net heat current across the system is zero, and consequently $\dot{Q}_{\text{c}}=-\dot{Q}_{\text{h}}$. Using the previous relation, we can express the heat current as
@@
where we used that $[\hat{J}_{+},\hat{J}_{-}]=2\hat{J}_{z}$. Furthermore, the case of a single qubit can be obtained from this equation setting $J=1/2$.
Notably, Eq. (23) reveals that the steady-state heat current does not explicitly depend on the sum of coherence terms such as $\langle\hat{J}_{+}\hat{J}_{-}\rangle$ and $\langle\hat{J}_{-}\hat{J}_{+}\rangle$, which are significant in the context of superradiant emission and is associated with the super-extensive scaling. Instead, the steady-state heat current is controlled by $\hat{J}_{z}$, which can be represented as a sum of $N$ local terms, $\hat{J}_{z}=\sum_{j=1}^{N}\hat{\sigma}^{(j)}_{z}$. This suggests the absence of super-extensive scaling due to superradiant effects in this context.
Surprisingly, despite this initial indication, as we will show below the steady-state heat current $\dot{Q}_{\text{h}}$ can exhibit a super-extensive scaling. This arises from the conservation of $\hat{J}^{2}$ and the consequent non-trivial occupation of different subspaces with different $J$. The steady-state density matrix $\hat{\rho}^{\rm(s)}$, as indicated by Eq. (17), is not purely thermal and encodes coherences between different qubits. As shown below, the presence of these coherences in this state can lead to a super-linear scaling of the heat current even at the steady state.
As a benchmark for our study, we will compare our results to the case where $N$ qubits (with equal frequency $\omega_{0}$) are independently coupled to a hot and cold bath. This scenario serves as our reference case, allowing for a clear comparison of our collective case described by a collective Lindblad master equation, Eq. (9). To compute the heat current associated with a single qubit, we employ the Lindblad master equation in Eq. (9), along with the formula for the steady-state heat current found in Eq. (23) in the case where $\bar{J}=1/2$, corresponding to $N=1$. As all qubits in our model are identical, the total heat current in this independent case, denoted as $\dot{Q}^{\rm ind}_{\rm h}$, is determined by multiplying the heat current of a single qubit by the total number of qubits, $N$. This leads us to the following expression for the total heat current in the independent scenario:
@@ II.3 Results for the heat current in a fixed subspace
For example, the case $\bar{J}=N/2$ can be implemented by initializing all the qubits in their ground states, |$\ket{0}_{j}$. The reason behind this is that the state $\bigotimes_{j=1}^{N}\ket{0}_{j}$ corresponds to the state |$\ket{N/2,-N/2}$, constraining the system’s dynamics to the subspace with $J=N/2$. While this simplified case may not be directly applicable in experimental settings, it serves as a useful starting point for gaining insights into this problem before we explore more realistic scenarios.
@@
In this particular setting, the steady-state heat current $\dot{Q}_{\rm h}$, as given in Eq. (24), can be analytically derived by averaging the current operator over the thermal state at an effective temperature denoted by $T^{*}$ and given in Eq. (16). The comprehensive calculation for this quantity is provided in Appendix C. Here, we only concentrate on discussing certain pertinent limits.
In the limit where the effective temperature $T^{*}$ is considerably lower than $T_{0}$ (specifically, $\bar{J}T_{0}\gg T^{*}$), the heat current can be approximated as follows:
@@
$\dot{Q}_{\text{h}}\approx\frac{2\gamma_{\text{h}}\gamma_{\text{c}}}{\gamma_{\text{h}}+\gamma_{\text{c}}}\hbar\omega_{0}\bar{J}\left(n_{\text{h}}-n_{\text{c}}\right)~{}.$ (27)
This expression can be interpreted by noticing that at low temperatures the majority of spins are in the inverted state, corresponding to $\langle\hat{J}_{z}\rangle\approx-\bar{J}$. Consequently, the heat current demonstrates a linear dependence on $\bar{J}$, a direct result of this spin inversion at low temperatures. Since $\bar{J}$ is at most linear in $N$, we do not have a super-extensive scaling of the heat current in this limit. In the opposite regime $\bar{J}T_{0}\ll T^{*}$, by means of Eq. (89), we have
$\dot{Q}_{\text{h}}\approx\frac{2\gamma_{\text{h}}\gamma_{\text{c}}}{\gamma_{\text{h}}+\gamma_{\text{c}}}\frac{(\hbar\omega_{0})^{2}}{3k_{\rm B}T^{*}}\bar{J}(1+\bar{J})\left(n_{\text{h}}-n_{\text{c}}\right)~{}.$ (28)
@@
In this scenario, the heat current exhibits a super-linear scaling with $\bar{J}$. This scaling behavior is derived from the non-linear dependence of the thermal populations $P(m_{\bar{J}}\,|\,\bar{J})$ on $\bar{J}$, given in Eq. (18), as shown in Appendix C. In the case $\bar{J}=N/2$, this corresponds to a super-extensive scaling in the number of qubits $N$. However, it is crucial to recognize that in the thermodynamic limit ($N\to\infty$), the condition $\bar{J}T_{0}\ll T^{*}$ is never met when $\bar{J}=N/2$. This is because, as $N$ approaches infinity, $\bar{J}=N/2$ eventually surpasses $T^{*}/T_{0}$.
@@
Figure 4: (Color online) Panels (a) shows the heat current as a function of the ratio $T_{0}/T_{\rm h}$ (where $T_{0}\equiv\hbar\omega_{0}/k_{\rm B}$) for various values of $\bar{J}$. The black dotted line indicates the independent current $Q_{\rm h}^{\rm ind}$, whereas the red dashed curve represents the heat current in the thermodynamic limit, denoted as $Q_{\rm h}^{\rm lim}$. In panel (b), the heat current is plotted as a function of $\bar{J}$ for various ratios of $T_{0}/T_{\rm h}$. The dashed lines in this panel correspond to the heat current in the thermodynamic limit, $Q_{\rm h}^{\rm lim}$, in the case $\bar{J}=N/2$. Each of these dashed lines corresponds to the same temperature ratio $T_{0}/T_{\rm h}$ of the data points of matching color. The temperature of the cold bath is chosen to be $T_{0}/T_{\rm c}=3$ in both panels. Other parameters are $\gamma_{\rm h}=\gamma_{\rm c}=\omega_{0}$.
Nonetheless, we still find a parametric advantage in the heat current. In the thermodynamic limit $N\to\infty$, we find that the ratio between the heat current $\dot{Q}^{\rm lim}_{\text{h}}$ in the $\bar{J}=N/2$ case, and the heat current in the independent case $\dot{Q}^{\rm ind}_{\text{h}}$, is given by
$\frac{\dot{Q}^{\rm lim}_{\text{h}}}{\dot{Q}^{\rm ind}_{\text{h}}}=\coth\left[\frac{T_{0}}{2T^{*}}\right]~{}.$ (29)
Although this result confirms that in the thermodynamic limit collective effects do not lead to a super-extensive scaling, the proportionality constant indicates a parametric advantage in certain scenarios. Specifically, in the case where $T_{0}/T^{*}\gg 1$, the collective heat current converges to the independent scenario, yielding a ratio ${\dot{Q}^{\rm lim}_{\text{h}}}/{\dot{Q}^{\rm ind}_{\text{h}}}\approx 1$. Conversely, in the limit where $T_{0}/T^{*}\ll 1$, the same ratio becomes ${\dot{Q}^{\rm lim}_{\text{h}}}/{\dot{Q}^{\rm ind}_{\text{h}}}\approx\left(2k_{\rm B}T^{*}/\hbar\omega_{0}\right)$, which is greater than one and unbounded. This indicates that, even in the thermodynamic limit, collective effects can provide an unbounded parametric enhancement to the heat current.
@@
In Fig. 4 (a,b) we plot the steady-state heat current $\dot{Q}_{\rm h}$ (normalized by $\bar{J}(\hbar\omega_{0})^{2}$) as a function of various parameters. This current has been calculated by finding the steady state of Eq. (14), by solving $\mathcal{D}(\rho^{\rm(s)})=0$. Fig. 4 (a,b) focuses on system sizes of experimental relevance, specifically considering a system size up to $N=16$, which corresponds to $\bar{J}=8$. For convenience, we fix the cold bath temperature $T_{c}=T_{0}/3$ and tune the ratio $T_{0}/T_{\rm h}$, which according to the discussion of the steady-state (see Fig. 3) allows tuning the effective temperature across the relevant scale $T_{0}$ as discussed above. In panel (a) we see that at high qubit temperatures, $T_{0}>1.5T_{\rm h}$ (corresponding to $T_{0}>T^{*}$ see Fig. 3), the ratio $\dot{Q}_{\rm h}/[\bar{J}(\hbar\omega_{0})^{2}]$ does not depend on $\bar{J}$, i.e. the heat current scales linearly in $\bar{J}$, thus in $N$. However, when the hot bath is hotter than $T_{0}$, i.e. $T_{0}<1.5T_{\rm h}$ (corresponding to $T_{0}<T^{*}$), a super-linear scaling emerges. In panel (a) we also show as a dotted line the heat current obtained in the thermodynamic limit $\dot{Q}_{\rm h}^{\rm lim}$ (taking $\bar{J}=N/2$) as given in Eq. (29). While a saturation to the extensive limit $\dot{Q}_{\rm h}^{\rm lim}$ is seen for low temperatures, there is no saturation for very high temperatures, $T_{0}/T_{h}\ll 1$, where the superlinear scaling persists up to $N=16$. The scaling with the size is explicitly shown in panel (b), where the dependence of the ratio $\dot{Q}_{\rm h}/[\bar{J}(\hbar\omega_{0})^{2}]$ on $\bar{J}$ is plotted for different temperatures, highlighting the superlinear scaling for small values of $T_{0}/T_{\rm h}$. Again, the heat current in the thermodynamic limit $\dot{Q}_{\rm h}^{\rm lim}$ is shown as dotted lines of correspondent color. Analogously to panel (a), if the hot bath has a temperature $T_{0}/T_{\rm h}\geq 0.7$ the heat current saturates the bound given by $\dot{Q}_{\rm h}^{\rm lim}$. Instead, for $T_{0}/T_{\rm h}\leq 0.3$ for realistic sizes the heat current always shows a superlinear behavior, far from saturation. To summarize, in these cases of practical experimental interest $N\leq 10$, if the hot bath is hot enough there is always a superlinear scaling and the value $\dot{Q}_{\rm h}^{\rm lim}$ rather represents an upper bound. Below we show that such a temperature regime that yields a superlinear scaling of the heat current is within experimental reach and that this phenomenology is resilient to noise.

### v3

@@ II.1 Steady-state density matrix and effective temperature
Since we are interested in computing the steady-state heat current flowing across the device, we start discussing the steady-state solution of the Lindblad master equation (9). In the case of a single reservoir, i.e. $\mathcal{D}(\hat{\rho})=\mathcal{D}_{i}(\hat{\rho})$ for $i=\text{h}$ or $i=\text{c}$, the steady-state solution corresponds to the thermal distribution (within each subspace with fixed $J$) characterized by the corresponding temperature $T_{i}$, since the rates satisfy the detailed balance condition
@@
capturing the equilibrium relation between absorption and emission rates in each bath. This result, consistent with the laws of thermodynamics, comes from using a master equation written in terms of the global jump operators for the coupled qubits (see Refs. 73, 74, 75 for a discussion of this point). In the case of two baths, the total dissipator in Eq. (10) can be expressed as:
@@
which, in turn, defines an effective temperature $T^{*}$, such that $T_{c}\leq T^{*}\leq T_{h}$. We can invert the previous equation to get an explicit expression for the effective temperature
@@
The steady-state solution $\hat{\rho}^{(\rm s)}$, determined setting $\mathcal{D}(\hat{\rho}^{(\rm s)})=0$, is thus a thermal state at temperature $T^{*}$ within each subspace with fixed $J$. However, since $\hat{J}^{2}$ is conserved, the relative occupation of different subspaces is fixed by the initial state. Assuming that the initial state does not contain any coherence between subspaces with different $J$, the stationary state $\hat{\rho}^{(\rm s)}$ is given by
@@
where $m_{J}$ are the eigenvalues of $\hat{J}_{z}$ compatible with $J$, $P_{J}$ is the occupation probability of subspace $J$ determined by the initial state, and |$\ket{J,m_{J}}$ are the Dicke states, given in Eq. (8). It should be noted that the summation over $J$ in Eq. (17) incorporates the degeneracies associated with $J$, as customary in angular momentum composition [76]. For example, considering a system composed of three qubits, each characterized by a spin of $\frac{1}{2}$, the resultant collective angular momentum can exhibit values of $J=1/2,1/2,3/2$, with the $J=1/2$ appearing twice. This is due to the composition rule that three spin-$\frac{1}{2}$ entities combine to yield collective angular momenta of $\frac{3}{2}$ and two instances of $\frac{1}{2}$. This composition rule is customarily represented as $\frac{1}{2}\otimes\frac{1}{2}\otimes\frac{1}{2}=\frac{3}{2}\oplus\frac{1}{2}\oplus\frac{1}{2}$. The dissipator in Eq. (14) fixes the occupation within each subspace $J$, i.e. it fixes the conditional probabilities $P(m_{J}|J)$ to be thermal at the effective temperature $T^{*}$:
@@
where $Z_{J,T^{*}}=\sum_{m_{J}}\exp\left[{-m_{J}T_{0}}/T^{*}\right]$ is the partition function.
@@ II.2 Heat Current
The general expression for the total heat current $\dot{Q}$ flowing out of the thermal baths can be defined as [77, 5]
@@
Using the specific forms of the dissipators for the cold and hot baths (in Eq. (11)), as detailed in App. C, we arrive at the following expressions for the heat currents
@@
where $\langle\hat{x}\rangle\equiv\text{Tr}[\hat{\rho}\hat{x}]$. This result is reminiscent of the well-known superradiant energy emission in an ensemble of excited qubits [37, 38, 78, 79]. In such a context, the system is typically considered to be in contact with a single cold bath at zero temperature, corresponding to setting $\gamma_{\rm h}=0$ in our case, and is initially prepared in a highly excited state, such as the state with $J=N/2$ and $M=J$. Consequently, the superradiant heat current flowing into the cold bath is given by $\dot{Q}_{\rm c}=-\hbar\omega_{0}\gamma_{\rm c}\langle\hat{J}_{+}\hat{J}_{-}\rangle$. The term $\langle\hat{J}_{+}\hat{J}_{-}\rangle$ can be expressed in terms of the individual qubit raising and lowering operators, $\hat{\sigma}_{+}^{(j)}$ and $\hat{\sigma}_{-}^{(j)}$, as $\sum_{j,j^{\prime}}\langle\hat{\sigma}^{(j)}_{+}\hat{\sigma}^{(j^{\prime})}_{-}\rangle$. This summation encompasses $N$ local population terms ($j=j^{\prime}$), and $\sim N^{2}$ non-local coherent terms ($j\neq j^{\prime}$) that may enable a super-extensive scaling effect. However, it is important to notice that this superradiant effect is only present in the transient dynamics. Over time, the qubits will eventually reach thermal equilibrium with the zero-temperature bath, resulting in the cessation of the heat current.
Going back to the case under study, our primary focus is on the steady-state heat current that flows between the two baths. Using that $\dot{Q}_{\text{c}}=-\dot{Q}_{\text{h}}$, we can express the heat current as
@@
where we used that $[\hat{J}_{+},\hat{J}_{-}]=2\hat{J}_{z}$. Notice that the case of a single qubit can be obtained from this equation setting $J=1/2$.
Notably, Eq. (24) reveals that the steady-state heat current does not explicitly depend on the sum of the non-local coherent terms that are present in $\langle\hat{J}_{+}\hat{J}_{-}\rangle$ and $\langle\hat{J}_{-}\hat{J}_{+}\rangle$, which are typically associated with a super-extensive behavior in the context of superradiant emission. Instead, the steady-state heat current is controlled by the operator $\hat{J}_{z}$, which can be represented as a sum of $N$ local terms, $\hat{J}_{z}=\sum_{j=1}^{N}\hat{\sigma}^{(j)}_{z}$. This would suggest the absence of a super-extensive scaling due to superradiant effects in this context. Indeed, upper bounding the expectation value $\langle-\hat{J}_{z}\rangle$ in Eq. (24) with $N/2$, we find the following exact upper bound
$\dot{Q}_{\text{h}}\leq\dot{Q}_{h}^{\text{max}}\equiv\frac{\gamma_{\text{h}}\gamma_{\text{c}}}{\gamma_{\text{h}}+\gamma_{\text{c}}}\hbar\omega_{0}N(n_{\text{h}}-n_{\text{c}})$ (25)
revealing that, for large values of $N$, the heat current cannot scale more than linearly in the number of qubit. It is interesting to notice that the bound in Eq. (25), which is saturable in the thermodynamic limit $N\to\infty$ (see Subsec. II.3), has a “Landauer-Büttiker-like” dependence on the temperature of the baths [80, 81], and is analogous to the heat current flowing through a harmonic oscillator coupled to bosonic heat baths [82] - but enhanced by a factor $N$. Intuitively this can be expected because, in the thermodynamic limit, the collective spin operators effectively become bosonic operators [83], and should thus behave similarly to a harmonic oscillator. However, the rates are here increased by a factor $N$ since the qubits are coupled in parallel.
Surprisingly, despite the linear scaling of $\dot{Q}_{\text{h}}^{\text{max}}$ with $N$, below we show that the heat current $\dot{Q}_{\text{h}}$ can nonetheless exhibit a super-extensive scaling in the finite-size regime. Indeed, if the qubit were in a global thermal state, it can be shown from Eq. (24) that the scaling would be linear, since the thermal state is a tensor product of local qubit states. However, since $\hat{J}^{2}$ is conserved, the steady-state density matrix $\hat{\rho}^{\rm(s)}$, as indicated by Eq. (17), may not be purely thermal, and the presence of coherence between different qubits can enable a non-linear dependence of the heat current on the qubit number $N$.
As a benchmark for our study, we will compare our results to the case where $N$ qubits (with equal frequency $\omega_{0}$) are independently coupled to a hot and cold bath. This scenario serves as a reference case, allowing for a clear comparison of our collective case described by a collective Lindblad master equation, Eq. (9). To compute the heat current associated with a single qubit, we employ the Lindblad master equation in Eq. (9), along with the expression of the steady-state heat current given in Eq. (24), by setting $\bar{J}=1/2$, corresponding to $N=1$. As all qubits in our model are identical, the total heat current in this independent case, denoted as $\dot{Q}^{\rm ind}_{\rm h}$, is determined by multiplying the heat current of a single qubit by the total number of qubits, $N$. This leads us to the following expression for the total heat current in the independent scenario:
@@
An analogous results for the $N=1$ was previously derived in the spin-boson model [82, 84].
@@ II.3 Results for the heat current in a fixed subspace
in Eq. (17). Notice that, in general, $\bar{J}\leq N/2$, and the case $\bar{J}=N/2$ can be implemented by initializing all the qubits in their ground states $\bigotimes_{j=1}^{N}\ket{0}_{j}$, corresponding to the Dicke state |$\ket{N/2,-N/2}$ and hence the system’s dynamics is constrained to the subspace with $J=N/2$. While this simplified case may not be directly applicable in experimental settings, it is a useful starting point for gaining insights into this problem before exploring more realistic scenarios.
@@
In this setting, the steady-state heat current $\dot{Q}_{\rm h}$, as given in Eq. (24), can be analytically derived by averaging the current operator over the thermal state at an effective temperature denoted by $T^{*}$ and given in Eq. (16). In the general case, the detailed calculation of this quantity is provided in Appendix C.
Analogously to the derivation of Eq. (25), we can upper bound the expectation value $\langle-\hat{J}_{z}\rangle$ in Eq. (24) with the conserved value of $\bar{J}$, yielding a generally tighter bound
$\dot{Q}_{\text{h}}\leq\dot{Q}_{\text{h}}^{\text{lim}}\equiv\frac{2\gamma_{\text{h}}\gamma_{\text{c}}}{\gamma_{\text{h}}+\gamma_{\text{c}}}\hbar\omega_{0}\bar{J}(n_{\text{h}}-n_{\text{c}}).$ (28)
Interestingly, as derived in Eq. (84), the heat current saturates this bound in the limit $\bar{J}T_{0}\gg T^{*}$. Physically, this can be interpreted as the low-temperature regime, i.e. when the effective temperature $T^{*}$ is considerably lower than $\bar{J}T_{0}$, or as the thermodynamic limit $N\to\infty$, choosing $\bar{J}=N/2$. Indeed, in these cases, the majority of spins are in the inverted state, corresponding to $\langle\hat{J}_{z}\rangle\approx-\bar{J}$. Since $\bar{J}$ is at most linear in $N$, this reveals that the heat current cannot scale super-extensively in the number of qubits both in the low-temperature regime and in the thermodynamic limit.
Nonetheless, in the thermodynamic limit $N\to\infty$ we still find a parametric advantage in the heat current. In this limit, the ratio between the heat current in the collective case $\dot{Q}_{\text{h}}^{\text{lim}}$ in the $\bar{J}=N/2$ case, and the heat current in the independent case $\dot{Q}^{\rm ind}_{\text{h}}$, is given by Eq. (86), i.e.
$\frac{\dot{Q}_{\text{h}}^{\text{lim}}}{\dot{Q}^{\rm ind}_{\text{h}}}=\coth\left[\frac{T_{0}}{2T^{*}}\right]~{}.$ (29)
The ratio ${\dot{Q}^{\rm lim}_{\text{h}}}/{\dot{Q}^{\rm ind}_{\text{h}}}$ is always greater than $1$, indicating that the collective scenario yields a strictly larger heat current than the independent case. In the limit $T_{0}/T^{*}\gg 1$, the ratio approaches one. Conversely, in the limit where $T_{0}/T^{*}\ll 1$, the ratio becomes ${\dot{Q}^{\rm lim}_{\text{h}}}/{\dot{Q}^{\rm ind}_{\text{h}}}\approx\left(2T^{*}/T_{0}\right)$, which can be arbitrarily large. This indicates that, even in the thermodynamic limit, collective effects can provide an unbounded parametric enhancement to the heat current.
We now consider the high-temperature regime, i.e. $\bar{J}T_{0}\ll T^{*}$. Using Eq. (85), we find that
@@
$\dot{Q}_{\text{h}}\approx\frac{2}{3}\frac{\gamma_{\text{h}}\gamma_{\text{c}}}{\gamma_{\text{h}}+\gamma_{\text{c}}}\frac{T_{0}}{T^{*}}(\hbar\omega_{0})\bar{J}(1+\bar{J})\left(n_{\text{h}}-n_{\text{c}}\right)~{}.$ (30)
@@
In this scenario, the heat current exhibits a super-linear scaling with $\bar{J}$. This scaling behavior is derived from the non-linear dependence of the thermal populations $P(m_{\bar{J}}\,|\,\bar{J})$ on $\bar{J}$, given in Eq. (18), as shown in Appendix C. In the case $\bar{J}=N/2$, this corresponds to a super-extensive scaling in the number of qubits $N$. However, it is crucial to recognize that in the thermodynamic limit ($N\to\infty$), the condition $\bar{J}T_{0}\ll T^{*}$ is never met when $\bar{J}=N/2$ since, as $N$ approaches infinity, $\bar{J}=N/2$ eventually surpasses $T^{*}/T_{0}$, and the upper limit $Q_{\text{h}}^{\text{lim}}$ in Eq. (28), linear in $\bar{J}$, becomes tight.
@@
Figure 3: (Color online) Panels (a) shows the heat current as a function of the ratio $T_{0}/T_{\rm h}$ (where $T_{0}\equiv\hbar\omega_{0}/k_{\rm B}$) for various values of $\bar{J}$. The black dashed line indicates the independent current $\dot{Q}_{\rm h}^{\rm ind}$, whereas the red dotted curve represents the heat current in the thermodynamic limit, denoted as $\dot{Q}_{\rm h}^{\rm lim}$. In panel (b), the heat current is plotted as a function of $\bar{J}$ for various ratios of $T_{0}/T_{\rm h}$. The dotted lines in this panel correspond to the heat current in the thermodynamic limit, $\dot{Q}_{\rm h}^{\rm lim}$, in the case $\bar{J}=N/2$. Each of these dotted lines corresponds to the same temperature ratio $T_{0}/T_{\rm h}$ of the data points of matching color. The temperature of the cold bath is chosen to be $T_{0}/T_{\rm c}=3$ in both panels. Other parameters are $\gamma_{\rm h}=\gamma_{\rm c}=\omega_{0}/10$.
@@
In Fig. 3 (a,b) we plot the steady-state heat current $\dot{Q}_{\rm h}$ (normalized by $\bar{J}\hbar\omega_{0}^{2}$) (14) focusing on system sizes of experimental relevance, specifically considering a system size up to $N=16$, which corresponds to $\bar{J}=8$. For convenience, we fix the temperature of the cold bath to be $T_{c}=T_{0}/3$ and tune the ratio $T_{0}/T_{\rm h}$, which according to the discussion on the steady-state’s effective temperature allows tuning the effective temperature across the relevant scale $T_{0}$ as discussed above. In panel (a) we see that in the low-temperature regime, $T_{0}>1.5T_{\rm h}$ (corresponding to $T_{0}>T^{*}$ ), the ratio $\dot{Q}_{\rm h}/(\bar{J}\hbar\omega_{0}^{2})$ does not depend on $\bar{J}$, i.e. the heat current scales linearly in $\bar{J}$, thus in $N$. However, in the high-temperature regime, i.e. $T_{0}<1.5T_{\rm h}$ (corresponding to $T_{0}<T^{*}$), a super-linear scaling emerges. In panel (a) we also show as a dotted line the heat current obtained in the thermodynamic limit $\dot{Q}_{\rm h}^{\rm lim}$ (taking $\bar{J}=N/2$), as given in Eq. (28), which is also an upper bound to the heat current. While, as expected, a saturation to the extensive limit $\dot{Q}_{\rm h}^{\rm lim}$ is seen for low-temperatures, there is no saturation for very high-temperatures, $T_{0}/T_{h}\ll 1$, where the superlinear scaling persists up to $N=16$. The scaling with the size is explicitly shown in panel (b), where the dependence of the ratio $\dot{Q}_{\rm h}/(\bar{J}\hbar\omega_{0}^{2})$ on $\bar{J}$ is plotted for different temperatures, highlighting the superlinear scaling for small values of $T_{0}/T_{\rm h}$. Again, the heat current in the thermodynamic limit $\dot{Q}_{\rm h}^{\rm lim}$ is shown as dotted lines of correspondent color. Analogously to panel (a), for low-temperatures $T_{0}/T_{\rm h}\geq 0.7$, the heat current saturates the bound given by $\dot{Q}_{\rm h}^{\rm lim}$. Instead, in the high-temperature regime, i.e. for $T_{0}/T_{\rm h}\leq 0.3$ and for realistic system sizes, the heat current always shows a superlinear behavior, far from saturation.
To summarize, for experimentally relevant cases where $N\leq 10$, a super-extensive scaling of the heat current is always achievable provided the hot bath is sufficiently warm (high-temperature regime). In these scenarios, the value of $\dot{Q}_{\rm h}^{\rm lim}$ essentially serves as an upper limit. Below we show that such a temperature regime that yields a super-extensive scaling of the heat current is within experimental reach and importantly, this behavior remains robust in the presence of noise.